OpenStudy (anonymous):

find the limit lim x approaches a x-1/sqrt. 2x=x^2-1

OpenStudy (anonymous):

\[\lim_{x \rightarrow a} \frac{ x-1 }{ \sqrt{2x-x^2-1} }\]

OpenStudy (goformit100):

Take Limit in use, with the fabonacci Summation

OpenStudy (anonymous):

a=1negative

OpenStudy (anonymous):

help

OpenStudy (goformit100):

ya correct.

sam (.sam.):

\[\large \lim_{x\to 1^{-}} \frac{(x-1) }{\sqrt{-x^2+2 x-1}}\]

OpenStudy (anonymous):

i work on this so far, whatt will be the next?\[\frac{ \sqrt{(x-1)^2} }{ -\sqrt{(x-1)(x-)}}\]

sam (.sam.):

Try factor it \[\frac{a-1}{\sqrt{-1}\sqrt{(a-1)^2}} \\ \\ \frac{a-1}{i(a-1)}\]

sam (.sam.):

i?

OpenStudy (anonymous):

ok thanks